# Autonomous Behavior and Whole-Brain Dynamics Emerge in Embodied Zebrafish Agents with Model-based Intrinsic Motivation

### v1

@@ 4 Results
We next assessed the ability of intrinsic motivation methods to replicate detailed behavioral patterns observed in biological zebrafish during autonomous exploration. Behavioral alignment was quantified by comparing locomotor trajectories and state transitions (active to passive and back) between artificial agents and zebrafish across multiple trials and subjects. 3M-Progress discovers ethologically-relevant behavioral state transitions by 10 million environment steps, whereas other intrinsic drives display transient strategies and stabilize on complete passivity or activity over training (Figure 3A). Agents trained with 3M-Progress exhibited the highest model-brain alignment, successfully capturing the dynamics of behavioral transitions recorded in the biological data (Figure 3B).
Additionally, agents trained with traditional intrinsic motivation methods such as ICM, RND, Disagreement, and $\gamma$-progress showed significantly lower alignment and failed to capture the characteristic stable cycling between activity states (Figure 3B-C).
@@
To evaluate how closely intrinsic motivation-driven agents matched zebrafish neural-glial dynamics, we compared model predictions to whole-brain calcium imaging data recorded during futility-induced passivity transitions. Neural-glial alignment was quantified using a One-to-One mapping, where each recorded biological neural or astrocytic unit was matched to the single most correlated artificial unit from the model. Strikingly, 3M-Progress agents captured nearly all of the explaining explainable variance in neural and astrocytic activity, markedly outperforming all existing intrinsic motivation methods, as well as baseline controls (Figure 4A). Model-free intrinsic drive controls include energy cost (homeostatic), entropy bonus (MaxEnt), and a randomly intialized agent. Data-derived controls include a Gaussian Process fit to neural-glial responses, proportional-integral-derivative (PID) control, average neural-glial population response, and white-noise. In fact, the 3M-Progress agent is the only model that is best aligned with both the behavioral alignment and neural-glial alignment, highlighting that accurate behavioral modeling closely corresponds with capturing detailed neural dynamics (Figure 4B), thereby passing the NeuroAI Turing test on this dataset [51].
@@
Figure 5: Latent dynamics of 3M-Progress agent’s internal activations compared with whole-brain neural-glial response in zebrafish. A) Principal components during passive and active transitions in the agent. B) Average whole-brain neural-glial response during passive and active transitions in a zebrafish subject.
Given that 3M-Progress best matched both behavioral and neural-glial alignment among all candidate models, we then characterized the internal dynamics of the 3M-Progress agent by performing Principal Component Analysis (PCA) on its in silico neural-glial population during active-to-passive state transitions. PCA revealed that the dominant latent dimensions of the agent’s core module closely mirrors whole-brain neural-glial dynamics measured from biological zebrafish, whereby glial responses accumulate evidence of motor futility via noradrenergic signaling to drive behavioral suppression, and neural responses reflect transient activation patterns associated with detecting mismatches between expected and actual sensory outcomes during unsuccessful swim attempts. This analysis demonstrates that the 3M-Progress intrinsic motivation mechanism not only generates realistic behavior, but also robustly captures fundamental internal neural-glial computations underlying autonomous exploration and behavioral state transitions (Figure 5).
@@ 5 Discussion
We introduced 3M-Progress, a novel intrinsic motivation formulation grounded in continuously tracking divergences between an internal world model and ethologically relevant memory, which robustly produced naturalistic transitions between active and passive behavioral states. Unlike prior intrinsic motivation methods that suffer from behavioral inconsistency and stochasticity, 3M-Progress reliably generated stable cycling behaviors closely matching those observed in biological zebrafish. Our work provides the first goal-driven computational model of neural-glial interactions and demonstrates its ability to replicate both behavioral patterns and underlying neural-glial computation observed during autonomous exploration in larval zebrafish.
Specifically, 3M-Progress agents were uniquely successful in capturing whole-brain neural-glial activity among all candidate models. We showed that artificial agent’s internal latent dynamics mirror neural-glial computation, wherein astrocytic responses accumulate evidence of motor futility through noradrenergic signaling to trigger behavioral suppression, while neural populations transiently encode mismatches between expected and observed sensory outcomes. Thus, our approach provides both behavioral, population-level, and mechanistic alignment with larval zebrafish.
Our findings suggest two complementary evolutionary constraints for developing robust, animal-like autonomous agents: (1) maintaining an intrinsic drive informed by memory, and (2) continuously monitoring divergences between this memory and the predictions of its world model. Functionally, tracking these mismatches within reinforcement learning frameworks allows artificial agents to identify ineffective strategies, update internal models adaptively, and redirect exploration towards more fruitful actions. Such mechanisms offer significant potential for enhancing the robustness and adaptivity of artificial autonomous systems, particularly in open-ended environments lacking clear external reward.
@@
Despite these strengths, several limitations merit consideration. Our current analyses primarily focused on passive behavioral transitions within constrained virtual environments, limiting ecological realism. Future work could extend both the artificial agents’ environments and the neural experiments to richer ecological settings and more complex behavioral repertoires, providing stronger and more diverse constraints on computational models of autonomy. Related to improving the richness of the environment, the biomechanical realism of the body could be further expanded to include muscles, providing additional constraints on the lower-level control problem. Additionally, while our model captures essential neuron-glial interactions at a population level, it abstracts away detailed biochemical signaling mechanisms within astrocytes and neurons. Incorporating more biologically detailed models of these subcellular processes could aid in providing predictions about signaling mechanisms at the subcellular scale. Nevertheless, our current approach provides a robust foundation for systematically investigating these open questions, paving the way for building more sophisticated autonomous artificial agents capable of adaptive exploration and flexible decision-making comparable to biological organisms.
@@ Acknowledgements
We thank Misha Ahrens, Chris Doyle, and Yu Mu for helpful discussions. A.N. was supported in part by a grant from the Burroughs Wellcome Fund. X.P. and R.K. were supported in part by the National Science Foundation and DoD OUSD (R & E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence, ARNI).
@@
In the descriptions below, we recycle several symbols between sections. Within sections, symbols are defined and self-consistent, so it should be clear to the reader what we are referring to. All experimental results involving policy rollouts are averaged across 8 random evaluation seeds.
Appendix A Model-Brain Alignment for Active Transitions
Figure 6: Noise-corrected Pearson’s $r$ correlation between whole-brain neural and glial units and artificial units from trained agents during active transitions.
Appendix B Model-Brain Alignment per Agent Module
Figure 7: Model-Brain Alignment for Passive Transitions per Module. Core and policy modules include each LSTM block along with hidden layers of their MLP decoders. The sensory module includes the last layer of the propcioceptive and vision encoders.
Appendix C Description of Control Models
@@
Maximum Entropy (MaxEnt) Agent
@@
$P(t)$ =$=\begin{cases}P_{\text{base}}(t),&G(t)\leq\theta_{G},\\ \\ 0,&G(t)>\theta_{G}.\end{cases}$
@@
Appendix D Implementation Details

### v2

@@ 5 Experimental Results
We first assessed the ability of intrinsic motivation methods to replicate detailed behavioral patterns observed in biological zebrafish during autonomous exploration. Behavioral alignment was quantified by comparing locomotor trajectories and state transitions (active to passive and back) between artificial agents and zebrafish across multiple trials and subjects.
Figure 5: Latent dynamics of 3M-Progress agent’s internal activations compared with normalized whole-brain neural-glial response in zebrafish. A) Principal components during passive and active transitions in the agent. B) Normalized average whole-brain neural-glial response during passive and active transitions in a zebrafish subject.
3M-Progress discovers ethologically-relevant behavioral state transitions by 10 million environment steps and sooner, whereas other intrinsic drives display transient strategies and stabilize on complete passivity or activity over training (Figure 3A). Agents trained with 3M-Progress exhibited the highest model-behavior alignment, successfully capturing the dynamics of state transitions recorded in the biological data (Figure 3B). Additionally, agents trained with traditional intrinsic motivation methods such as ICM, RND, Disagreement, and $\gamma$-progress showed significantly lower alignment and failed to capture the characteristic stable cycling between active and passive states (Figure 3B-C).
@@
To evaluate how closely intrinsic motivation-driven agents matched zebrafish neural-glial dynamics, we compared model predictions to whole-brain calcium imaging data recorded during behavioral transitions of futility-induced passivity. Neural-glial alignment was quantified using a One-to-One mapping, where each recorded biological neuron and astrocyte was matched to the single most correlated artificial unit from the model. Strikingly, 3M-Progress agents captured nearly all of the explainable variance in neural and astrocytic activity, markedly outperforming existing intrinsic motivation algorithms, as well as baseline controls (Figure 4A). Model-free intrinsic drive controls include energy cost (homeostatic), entropy bonus (MaxEnt), and a randomly intialized agent. Data-derived controls include a Gaussian Process fit to neural-glial responses, proportional-integral-derivative (PID) control, average neural-glial population response, and white-noise. More details on these controls can be found in Appendix D. In fact, the 3M-Progress agent is the only model that is almost completely aligned with both the behavioral alignment and neural-glial alignment, highlighting that accurate behavioral modeling can tightly constrain detailed neural dynamics (Figure 4B). Taken together, 3M-Progress agents pass the NeuroAI Turing Test on this dataset, a criterion emphasizing models that match both behavior and internal function [65].
@@
Given that 3M-Progress best matched both behavioral and neural-glial alignment among all candidate models, we then characterized the agent’s internal dynamics by performing Principal Component Analysis (PCA) on its in silico neural-glial population during behavioral state transitions. PCA revealed that the dominant latent dimensions of the agent’s core module closely mirrors whole-brain neural-glial dynamics measured from biological zebrafish, whereby glial responses accumulate evidence of motor futility via noradrenergic signaling to drive behavioral suppression, and neural responses reflect transient activation patterns associated with detecting mismatches between expected and actual sensory outcomes during unsuccessful swim attempts. This analysis demonstrates that the 3M-Progress mechanism for intrinsic motivation not only generates realistic behavior, but also robustly captures fundamental internal neural-glial computations underlying autonomous exploration and behavioral state transitions (Figure 5).
@@ 6 Discussion
Our work seeks to identify and computationalize intrinsic goals that enable autonomous, task-independent behavior in animals. Leveraging a unique whole-brain dataset recorded in larval zebrafish during an autonomous behavior known as futility-induced passivity, we identify two simple principles of intrinsic goals for autonomous agents: avoid perseveration on stimuli that are uncontrollable and unpredictable, and converge on a robust decision-making strategy. We introduced 3M-Progress, a novel intrinsic drive that operationalizes these principles by continually tracking divergence between an online world model and an ethologically relevant prior. Learned from experience in an ecological niche that captures the basic physics of a naturalistic zebrafish habitat, this prior guides exploration in new environments by partitioning the behavioral space into niche-seeking and niche-avoiding modes.
Unlike prior intrinsic motivation methods that suffer from behavioral inconsistency and non-stationarity, 3M-Progress reliably generated stable cycling behaviors closely matching those observed in biological zebrafish. Moreover, 3M-Progress agents were uniquely successful in capturing whole-brain neural-glial activity among all candidate models. We showed that artificial agent’s internal latent dynamics mirror neural-glial computation, wherein astrocytic responses accumulate evidence of motor futility through noradrenergic signaling to trigger behavioral suppression, while neural populations transiently encode mismatches between expected and observed sensory outcomes. To the best of our knowledge, our work marks the first goal-driven model of neural-glial computation, as well as the first completely self-supervised embodied agent that predicts behavioral and brain data.
Our findings suggest two complementary evolutionary constraints for developing robust, animal-like autonomous agents: (1) maintaining an intrinsic drive informed by memory, and (2) continually monitoring divergence between this memory and new sensory experiences. Functionally, tracking this mismatch within reinforcement learning frameworks allows artificial agents to identify ineffective strategies, update internal models adaptively, and discover new behaviors. Such mechanisms offer significant potential for enhancing the autonomy of artificial systems, especially in open-ended environments lacking clear external rewards or goals.
@@
Our current analyses primarily focused on behavioral transitions within constrained virtual environments, limiting ecological realism. Future work could extend both the artificial environments and biological experiments to richer ecological settings and more complex behavioral repertoires, providing stronger and more diverse constraints on computational models of autonomy. The biomechanical realism of the body could also be expanded to include muscles and motor circuits, providing realistic constraints on the low-level controller. Additionally, while our model captures essential neuron-glial interactions at a population level, it abstracts away detailed biochemical signaling mechanisms and anatomy of astrocytes and neurons. Incorporating more biologically detailed models of these processes could aid in providing predictions about these mechanisms at a finer scale. Finally, 3M-Progress can be generalized in a variety of ways that extend beyond futility-induced passivity (see Appendix C.3)—we leave this to future work.
@@ Acknowledgements
We thank Misha Ahrens, Chris Doyle, and Yu Mu for helpful discussions, as well as the anonymous reviewers for their helpful feedback on the initial manuscript draft. A.N. was supported in part by a grant from the Burroughs Wellcome Fund. X.P. and R.K. were supported in part by the National Science Foundation and DoD OUSD (R & E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence, ARNI) and the Simons Collaboration in Ecological Neuroscience (SFI-AN-NC-SCN-00007276-10).
@@
Appendix A Model-Brain Alignment per Transition and per Module
Figure 6: Model-Brain alignment for active and passive transitions and per module, excluding per module read-in and readout layers where applicable. A) Alignment for Active Transitions. B) Alignment per Agent Module for Active Transitions. C) Alignment for Passive Transitions. D) Alignment per Agent Network Module for Passive Transitions.
Appendix B Latent Dynamics of Baseline Agents
Figure 7: Latent dynamics of each agent-based control model. Dashed line indicates time since passivity. Coloring of PCs neural-glial cell-types (Fig. 2, Fig. 5) was chosen according to cell-types of the 3M-Progress agent PCs.
Appendix C Formal Intuitions
C.1 Curiosity-driven Exploration is Not Enough
Consider an MDP with discount $\gamma\in(0,1)$, occupancy $d_{\pi}(s,a)=(1-\gamma)\sum_{t\geq 0}\gamma^{t}\Pr_{\pi}(s_{t}{=}s,a_{t}{=}a)$, and transition kernel $\mathcal{T}(s^{\prime}\mid s,a)$. Let a predictive world model with parameters $\theta$ be trained online to minimize
$\mathcal{L}(\pi,\theta)\;=\;\mathbb{E}_{(s,a,s^{\prime})\sim d_{\pi}\mathcal{T}}\big[\ell(s,a,s^{\prime};\theta)\big],$
and suppose the policy receives intrinsic reward determined by this predictor: $r^{i}(s,a,s^{\prime};\theta)\;=\;g\big(\ell(s,a,s^{\prime};\theta)\big)$ where $\ell\geq 0$ and $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is monotone increasing with $g(0)=0$. The policy is updated to maximize
$J(\pi,\theta)\;=\;\mathbb{E}_{(s,a,s^{\prime})\sim d_{\pi}\mathcal{T}}\!\big[r^{i}(s,a,s^{\prime};\theta)\big].$
Under either alternating or simultaneous updates that decrease $\mathcal{L}$ in $\theta$ and increase $J$ in $\pi$, the process cannot converge to a nontrivial, uniquely defined exploratory policy. The only stationary outcome is a degenerate collapse in which $r^{i}$ is constant on the visited support; otherwise the policy-model coupling remains non-stationary and induces drift. That is, any stationary point $(\pi^{\star},\theta^{\star})$ satisfies one of the following cases:
  1. 1.
    Reward collapse. If $\theta^{\star}$ minimizes $\mathcal{L}(\pi^{\star},\theta)$ on the support of $d_{\pi^{\star}}\mathcal{T}$ and the model class is realizable on that support, then $\ell(\cdot;\theta^{\star})=0$ almost surely, hence $r^{i}(\cdot;\theta^{\star})=0$ almost surely and $J(\pi,\theta^{\star})=0$ for all $\pi$. The objective is flat and does not select a unique exploratory policy.
  2. 2.
    No stationary policy. If we freeze the model at $\theta^{\star}$, then $r^{i}(\cdot;\theta^{\star})$ is a fixed reward. If residual error remains, then $\ell(\cdot;\theta^{\star})$ (hence $r^{{i}}$) is not almost surely constant, so there exists a measurable set $U$ on which the intrinsic advantage $A^{r^{\mathrm{int}}}_{\pi^{\star}}(s,a)>0$ for some actions. By the policy-gradient identity,
    $\nabla J(\pi^{\star},\theta^{\star})\;=\;(1-\gamma)\,\mathbb{E}_{d_{\pi^{\star}}}\!\big[A_{\pi^{\star}}(s,a)\,\nabla\log\pi^{\star}(a\mid s)\big].$
    For some state $s\in U$, suppose we increase $\pi^{\star}(a\mid s)$ slightly on an action with $A_{\pi^{\star}}(s,a)>0$ and decrease it on other actions to preserve $\sum_{a}\pi^{\star}(a\mid s)=1$. This choice makes the inner product $\langle A_{\pi^{\star}}(s,\cdot),\nabla\log\pi^{\star}(\cdot\mid s)\rangle$ strictly positive on $U$, hence the expectation above is positive and the policy gradient is nonzero. Therefore $\pi^{\star}$ is not a local maximizer of the stationary objective. Any such occupancy shift then triggers predictor updates that reduce $r^{\mathrm{int}}$ on $U$, moving the high-reward region and preventing a nontrivial fixed point.
In both cases, these intrinsic signals do not converge to a stable, uniquely-defined exploratory policy. Either training drives the intrinsic signal to a constant (degenerate) value on the visited distribution, or residual heterogeneity in the reward keeps creating ascent directions that are then neutralized by predictor updates, preventing stabilization.
C.2 3M-Progress Partitions the Behavioral Space
Consider two reward-free MDPs, $\mathcal{M}_{1}=\left({\mathcal{S}},{\mathcal{A}},\mathcal{T}_{1},{p}_{0}\right)$ and ${\mathcal{M}}_{2}=\left({\mathcal{S}},{\mathcal{A}},{\mathcal{T}}_{2},{p}_{0}\right)$, that differ only in their transition densities. Further, suppose there exists a set $U\subset\mathcal{S}\times\mathcal{A}$ such that $\mathcal{T}_{1}(\cdot\mid s,a)=\mathcal{T}_{2}(\cdot\mid s,a)$ almost everywhere in $s^{\prime}$ for all $(s,a)\in U$. Let $p(\cdot\mid s,a)$ and $q(\cdot\mid s,a)$ be world models trained on data from $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, respectively, with pointwise consistency:
$\mathcal{D}_{\mathrm{KL}}\left(\mathcal{T}_{1}(\cdot\mid s,a)\|p(\cdot\mid s,a)\right)\overset{p}{\rightarrow}0,\quad\mathcal{D}_{\mathrm{KL}}\left(\mathcal{T}_{2}(\cdot\mid s,a)\|q(\cdot\mid s,a)\right)\overset{p}{\rightarrow}0.$
Assume for each $(s,a)$, both kernels share the same support with no vanishing probabilities (i.e. $\exists c>0$ s.t. $\mathcal{T}_{2}(s^{\prime}\mid s,a)\geq c$ whenever $\mathcal{T}_{1}(s^{\prime}\mid s,a)>0)$. Then it follows that for all $(s,a)\in U$,
$\mathcal{D}_{\mathrm{KL}}\left(p(\cdot\mid s,a)\|q(\cdot\mid s,a)\right)\overset{p}{\rightarrow}0,$
and that for all $(s,a)\in U^{c}$,
$\mathcal{D}_{\mathrm{KL}}\left(p(\cdot\mid s,a)\|q(\cdot\mid s,a)\right)\overset{p}{\rightarrow}\mathcal{D}_{\mathrm{KL}}(\mathcal{T}_{1}\|\mathcal{T}_{2})>0.$
C.3 Beyond Futility-induced Passivity
Although we demonstrate 3M-Progress on a specific autonomous behavior known as futility-induced passivity, our algorithm applies to any exploration behavior in which a dynamics niche is reasonably specified. In our experiments, futility-induced passivity arises completely from the choice of the pretraining environment and the online learning environment; we choose these environments such that the transition dynamics between environments agree when the agent is passive, thus defining the ecological niche that guides exploration (see Appendix C.2: simply put, we choose environments such that there exists a subset $U$ in which the transition dynamics locally agree). In general, the pretraining environment should encode a meaningful dynamics prior the agent can use for continual learning with environments whose physics vary systematically from the pretraining environment.
For example, suppose we pretrain an agent and world model on a foraging task (e.g. the virtual rodent environment in [66]). In a new environment that includes the opportunity for foraging and any of its constituent locomotor primitives (running, turning, jumping, etc.), the basic distribution-matching objective of 3M-Progress rewards the agent for trajectories whose dynamics are predictable under the pretrained world model (namely, foraging and constituent motor primitives):
$r_{t}^{i}=f(\hat{\epsilon}_{t}-\epsilon_{t});\quad\epsilon_{t}=\mathcal{D}_{\mathrm{KL}}(\omega_{\theta^{\prime}}\|\omega_{\theta});\quad\hat{\epsilon}_{t}=(1-\gamma)\hat{\epsilon}_{t}+\gamma\epsilon_{t}$
where $\theta^{\prime}$ is learned online and $\theta$ is pretrained. Niche-aware exploration is primarily mediated by the activation function $f$. When $f$ is monotonic non-decreasing, such as a rectified linear unit, 3M-Progress is niche-seeking with the niche defined by the dynamics prior $\omega_{\theta}$. Conversely, if $f$ is monotonic non-increasing, 3M-Progress is niche-avoiding and explores dynamics outside the prior. Non-monotonic functions allow some amount of symmetry between niche-seeking and niche-avoidance depending on the specific function shape.
The utility of this approach can also be appreciated when the pretraining stage involves a large diversity of dynamics, either from multiple environments and tasks or a single environment with multiple tasks. A world model with sufficient computational capacity that captures these diverse dynamics can be flexibly used in new environments to motivate exploration in a variety of ways. In the simplest case, for example, suppose we pretrain an ensemble of dynamics models $\{\omega_{\theta_{j}}\}_{j=1}^{N}$ on $N$ separate environments and tasks. Maintaining independent temporal filters $\hat{\epsilon}_{t}^{j}$ for each prior, a deterministic intrinsic motivation can be defined as $r_{t}^{i}=\max\{f(\hat{\epsilon}_{t}^{j}-\epsilon_{t}^{j})\}_{j=1}^{N}$. Alternatively, one can imagine various sampling schemes over the ensemble in order to drive specific exploration styles, such as $\epsilon$-greedy or max-entropy. This extends 3M-Progress to niche-aware exploration over multiple niches, thereby allowing the flexibility of multiple modes of behavior in a single objective function. Each addition of a dynamics prior further partitions the behavioral space in the online environment, and allows the dynamics characteristic of each pretraining environment to be composed to form an exploration landscape of attractors (niche-seeking) or repellors (niche-avoidance) for online learning.
Appendix D Description of Control Models
D.1 Model-based Controls
The Intrinsic Curiosity Module (ICM)
[3] defines intrinsic reward as the Shannon surprise of the forward model, $r_{t}^{i}:=-\log\omega_{\theta}(\phi^{I}_{t}\mid\phi_{t},\mathbf{a}_{t})$. $I$ denotes an augmented inverse feature space that is learned on-top of sensory embeddings from $\phi$—ICM trains an additional embedding layer $\phi^{I}$ using an inverse dynamics model parameterized by $\theta_{I}$ by optimizing an MSE loss $L(\theta_{I})=\mathbb{E}_{\pi_{\theta}}\left\|f(\theta_{I};\phi^{I}_{t},\phi^{I}_{t+1})-\mathbf{a}_{t}\right\|_{2}^{2}$.
Random Network Distillation (RND)
[4] defines a fixed random nonlinear projection of sensory features $g(\phi)$ and trains a predictor network $\hat{g}$ using an MSE loss $r_{t}^{i}:=L(\theta_{\mathrm{RND}})=\mathbb{E}_{\pi_{\theta}}\left\|g(\phi_{t})-\hat{g}(\theta_{\mathrm{RND}};\phi_{t})\right\|_{2}^{2}$. With the distillation objective as the intrinsic reward, RND does not reinforce behaviors by scoring their predictability by a forward dynamics model as in ICM; instead, the random memory provides a simple exploration bonus for visiting novel states under the policy distribution.
Disagreement
[5] learn an ensemble of world models $\{\omega_{\theta_{j}}\}_{j=1}^{N}$ and defines intrinsic reward as $r_{i}^{t}:=\mathrm{Var}\left(\{\mu_{\theta_{j}}:j\in[N]\}\right)$ for $N$ randomly initialized world models. Ensemble variances scores the stochasticity of the environment and reinforces state-action pairs for which models disagree.
$\gamma$-Progress
[7] leverages the temporal history of Shannon surprise to define intrinsic reward using prediction gain, $r_{i}^{t}:=\log\frac{\omega_{\theta_{new}}}{{\omega_{\theta_{old}}}}$, where $\theta_{new}$ parameterized a world model after learning on new transitions withheld from an lagging model $\theta_{old}$.
D.2 Model-Free Controls
@@
Maximum Entropy Agent
@@
$P(t)$ $=\begin{cases}P_{\text{base}}(t),&G(t)\leq\theta_{G},\\ \\ 0,&G(t)>\theta_{G}.\end{cases}$
@@
Appendix E Implementation Details